OpenStudy (anonymous):

Jessica wants to record her favorite songs to one CD. The function C=80-3.22n represents the recording time C available after n songs are recorded. Find the zero of this function. Describe what this value means in this context.

OpenStudy (anonymous):

We put y and x as zeros to get the zero of a function. Lets start with x. C = 80 - 3.22n C = 80 - 3.22(0) C = 80 - 0 C = 80 So C = 80, when x = 0 Now we do y (C is equal to the y axis when on the graph, because that is the outcome of the manipulation) We do inverse operation. C = 80 - 3.22n 0 = 80 - 3.22n (subtract) -80 = 3.22n (divide) -24.8447 = n When C = 0, n = -24.8447

OpenStudy (anonymous):

thanks